### v1

@@ VI.3 Experimental progress on
Since SF II, five new experiments have been reported: Four from the Madrid and ATOMKI groups by the activation technique, at relatively high center of mass energies [2012PhRvC..86c2801C, 2013NuPhA.908....1B, 2019PhRvC..99e5804S, 2023PhRvC.108b5802T], and one from the Notre Dame group using the prompt- method [2013PhRvC..87f5804K].
@@
Following the approach adopted in SF II, we again only use the recoil and activation data for the fits below. The -ray angular distribution is not known experimentally [2007PhRvC..76e5801B], and the resultant uncertainty increases the common-mode error for the prompt- studies somewhat. (A related experiment has recently concluded at Felsenkeller Dresden. The data suggest a higher than expected anisotropy but are so far only available in the form of a PhD thesis [Turkat23-PhD].) In addition, all of the “modern” works reporting prompt- data except for Notre Dame [2013PhRvC..87f5804K] also include data obtained with other methods, leading to partial correlations between data sets that would complicate fitting.
@@
Following this restriction, our data selection proceeds as follows. First, the four data sets previously used in SF II are carried over here: Weizmann [2004PhRvL..93z2503S], LUNA [2006PhRvL..97l2502B, 2007PhRvC..75f5803C, 2007PhRvC..75c5805G] (only the activation data), Seattle [2007PhRvC..76e5801B] (only the activation data), and ERNA [2009PhRvL.102w2502D].
@@
Two of the archival data sets excluded by the 1998 cutoff date in SF II merit some further discussion here: The data of 1982PhRvL..48.1664O, 1984NuPhA.419..115O consist of two points measured by activation in a 3He gas cell. However, their uncertainties are not separated into statistical and systematic components, making them unsuitable for the data model used below. The work by 1983PhRvC..27...11R consists of one data point that was obtained by averaging two separate activation measurements with a 3He and a 4He gas cell, respectively. We do not use this result because we lack details to verify the background subtraction. We also choose not to use the activation study by 1983ZPhyA.310...91V. In that experiment there were thick entrance foils and the 4He beam was completely stopped inside a high-pressure 3He gas cell. This gave an integrated measurement over a wide energy range, so that the analysis depends strongly on the assumed shape of the curve, and the result was reported only as an extrapolated .
@@
We now consider the new data since SF II. An activation experiment in Madrid [2012PhRvC..86c2801C] used a 3He beam incident on a 4He gas cell, reporting three data points. Another activation experiment was reported by the ATOMKI group, using a 4He beam on 3He gas cells [2013NuPhA.908....1B]. Two higher-energy campaigns at ATOMKI were again performed using the activation method. These latter data are at = 2.5-4.4 MeV [2019PhRvC..99e5804S] and = 4.3-8.3 MeV [2023PhRvC.108b5802T], respectively, above the energy range suitable for halo EFT, and therefore not included.
@@
A detailed study by the Notre Dame group reported 17 data points [2013PhRvC..87f5804K], using the primary -ray from ground state capture and the secondary -ray from the deexcitation of the 429 keV first excited state of 7Be. No activation data are reported in that work, and the -ray detector was placed at just one angle, 90∘. Since only prompt- data are reported, we did not include this data set. In order to test the effects of this decision, we repeated some of our fits (see below, Section VI.4) with the modern prompt- experiments included: the 17 points from 2013PhRvC..87f5804K, 3 prompt- points from LUNA [2007PhRvC..75f5803C] and 8 from Seattle [2007PhRvC..76e5801B]. The extrapolated -factor changed by less than 1%, well within the error bars for the recommended .
@@
Finally, an indirect experiment using the 6Li(3He,)7Be reaction and the ANC technique has recently been reported [2020PhLB..80735606K]. It is left out of our fits due to the additional normalization and theory uncertainties involved in determining an ANC from a transfer experiment, which are larger than the errors in the factor measurements used for the present fits. All the data used in the present fits are summarized in Table 4.
@@
As in SF-II, we model data uncertainties as consisting of a component that is independent for each point and a common-mode component that applies to all data from a given experiment as a multiplicative factor. This separation is well-documented for all of the modern data and is also shown in Table 4. Except at the lowest energies, the common-mode error typically dominates.
@@ VI.4 Data fitting
We carried out both frequentist fits that minimize and Bayesian fits based on the posterior probability distribution of a likelihood function computed from . In the Bayesian fits parameters, extrapolated , and their errors were determined by Monte Carlo sampling of the parameter space. The relatively large number of model parameters and the significant fitting degeneracies between some of them make Bayesian analysis a natural choice for finding best fit and confidence intervals for the multi-parameter models applied here. Two of the groups did not produce frequentist fits in which they had confidence.
@@
One set of fits for the 2018EPJA...54...89H halo EFT was produced by minimization, proceeding in two steps: first LO parameters were fitted to keV data, and then the results were taken as initial values in the search to minimize over all parameters in the NLO theory for the full set of data. These fits gave keV b with the scattering constraint and keV b without (uncertainties being propagated in linear approximation using covariances and partial derivatives). Formally the EFT truncation error from stopping at NLO corresponds to an additional theoretical error that can shift factors by 10%. However, any fitted curve is constrained by low-energy data, which in some sense become effectively renormalization conditions of the field theory; the error on extrapolated should probably be smaller than 10% by an amount that is hard to estimate.
@@ VI.5 Recommended value
Because the fit to the NLO EFT of 2018EPJA...54...89H gives the central result among those attempted here, we adopt it for our recommended . For ease of adoption, it is noted that this recommended curve can be empirically parameterized using the same shape as in SF II, by the numerical equation (following a customary form in the past literature)
@@
Propagation of uncertainties from measurement to extrapolated is relatively straightforward and unambiguous (at least in the Bayesian fitting). However, the uncertainty due to theoretical approximation is more complicated, and error estimation on extrapolated quantities in EFTs remains an open area of research; estimates from the groups working on are discussed above. We estimate the overall theoretical uncertainty on our recommended factor to be the measured by the 4% dispersion among fitted curves; this is comparable to the more formal estimates.
Our recommendations for the zero energy astrophysical factor and the low-energy slope are then
@@
The first error bar for is given by the fit uncertainty, conservatively using the largest value among the various fits, and it encodes the experimental uncertainty. The second error bar reflects the 4% dispersion among theoretical formalisms. The two errors can be added in quadrature for applications. Our recommendation is essentially the same as in SF II, but derived from completely different theoretical approaches that are much more directly connected to the data and much less dependent on model assumptions. It is also based on data over a wider energy range. The dispersion of and among models at the solar Gamow peak is illustrated in Table 5, along with some indication of how fitted scattering lengths and squared ANCs vary among frameworks. (The latter are highly degenerate with short-range parameters in the fitting; in the 2018EPJA...54...89H EFT they are sensitive to small variations in the -wave effective ranges that are fitted instead of ANCs.)
@@ VI.6 Recommendations for future work on
Our third recommendation is a study of the astrophysical factor over a wide energy range, 50-1500 keV and extending to even lower energies when technically feasible, with small point-to-point errors. Even with a large common-mode error, this would significantly constrain theoretical curve shapes near threshold, and providing an independent check on the influence of scattering data on .
@@
VII The reaction ()
The reaction , also known as the hep reaction, is one of the possible processes taking place after the , and it produces the most energetic neutrinos, with an endpoint energy of 18.8 MeV. It is however the least probable, with its factor, and consequently its rate, even beyond the reach of current experiments. This is due to the fact that the hep reaction is induced by the weak interaction and further suppressed by Coulomb barrier and by the fact that the leading one-body axial (Gamow-Teller) current operator cannot connect the main -state components of the and initial- and final-state wave functions. As a consequence, the one-body axial current non-zero contribution is due to the small components of the and wave functions. Furthermore, other contributions, such as weak magnetism and other one-body corrections to the vector current, and two-body vector and axial currents, and the -waves in the initial state, normally suppressed at solar temperatures, are significantly enhanced. This is further complicated by cancellations between the one-body and two-body axial current contributions.
Due to this large suppression of the contributions that are nominally expected to be the largest, the zero-energy factor, , is unreachable by experiment and only theory can provide an estimate. The most recent studies of this reaction, already evaluated in SF-II, are those of 2000PhRvC..63a5801M and 2003PhRvC..67e5206P. The first one was performed within a phenomenological approach, similar to the calculation of 1998PhRvC..58.1263S for the pp fusion. In this particular case, the and initial and final nuclear wave functions were obtained with the correlated Hyperspherical Harmonics (CHH) variational method [1995FBS....18...25V, 1998PhRvL..81.1580V], using the Argonne (AV18) two-nucleon potential [1995PhRvC..51...38W] augmented by the Urbana IX (UIX) three-nucleon interaction [1995PhRvL..74.4396P]. In this way, the binding energies of 3He and 4He, and the singlet and triplet scattering lengths were in good agreement with experiment. The weak vector and axial transition operators were obtained within a phenomonological approach. In particular, the largest two-body contribution to the axial current arising from the excitation of intermediate -isobar degrees of freedom was included using the transition-correlation operator scheme of 1991PhRvC..44..619C, 1992PhRvC..45.2628S and fixing the nucleon-to- axial coupling constant to reproduce as in the pp case [1998PhRvC..58.1263S].
@@ VII The reaction ()
Following the results of 2000PhRvC..63a5801M and 2003PhRvC..67e5206P, SF-II recommended for the value keV b. Three recommendations were also made: (i) to perform new studies with a broad spectrum of Hamiltonian models, in order to properly access the theoretical uncertainty; (ii) to study other weak reactions for which experiments can test the predictions made by employing the same theoretical ingredients as those used for the hep reaction calculation; (iii) to further understand the relation between the hep reaction and the so-called hen process (). Of these recommended works, only the second one has been performed so far, applying the theoretical framework used for the hep study to the muon capture reactions on light nuclei [2011PhRvC..83a4002M].
Given the lack of further studies on the hep reaction after SF-II, we decide to maintain
as recommended value. Notice that we could have increased by %, in analogy with , due to the updated values of input parameters (mostly ). However, we have decided not to apply such an increase, because it does not derive from a direct calculation, and, most of all, because such an increase lies well within the quoted % uncertainty. However, given the advances in ab initio studies of the last decade and in the development of the EFT framework, we update the recommendations of SF-II as follows: (i) the hep process should be revisited within a fully consistent EFT approach, similarly to what has been done in these years for the pp fusion and other processes, as, for instance, muon captures on light nuclei [2012PhRvL.108e2502M, 2018PhRvL.121d9901M, 2018PhRvC..98f5506A, 2023FrP....1049919C, 2023PhRvC.107f5502B, PhysRevC.109.035502]; (ii) the initial and final wave functions should be calculated with the most accurate ab initio methods, as the (uncorrelated) HH method [2008JPhG...35f3101K, 2020FrP.....8...69M], which has been proven to provide more accurate bound and scattering wave functions than the CHH method, using both local and non-local interactions; (iii) the relation between the hep and the hen process should be understood. The hen process has been studied within EFT and the HH method by 2022PhRvC.105a4001V, but a consistent parallel study of both reactions is still missing.
@@
VIII Electron capture by and
Electron capture reactions are the sources of lines found in the solar neutrino spectrum. In particular, electron capture on , i.e., the process , also known as the pep reaction, competes with fusion and depends on the same nuclear matrix element. Therefore, the ratio between the pp and pep rates is independent of nuclear physics. Based on this consideration, in SF-II the result of SF-I, which is based on the work of 1969ApJ...155..501B, was multiplied by the radiative corrections calculated by 2003PhRvC..67c5502K, leading to the final result given by
where is the density in units of g/cm3, is the mean molecular weight per free electron, and is the temperature in units of MK. The range of validity is . Given that no new evaluation of the pep rate has been performed since SF-II, the result of Eq. (23) represents also the present recommended value.
Competition between electron and proton capture on 7Be fixes the branching ratio of the ppII and ppIII chains, and thereby the 7Be and 8B neutrino fluxes. Superallowed electron capture on the ground state of 7Be leads to the ground state or the , 478 keV first excited state of 7Li with a measured (terrestrial) branching ratio of 10.44(4)% [2002NuPhA.708....3T]. It is customary to calculate the solar decay rate in terms of the measured terrestrial decay rate, which has the advantage of removing the nuclear physics dependence, but it requires a proper calculation of the densities of electrons at the nucleus. This has usually been done by considering continuum and bound electrons separately, using the Debye-Hückel (DH) approximation pioneered by 1954AuJPh...7..373S for the screening of bound electrons in the solar plasma; the capture rate of continuum electrons is not significantly influenced by electron screening. Furthermore, it is necessary to calculate the atomic probability densities governing the and terrestrial electron capture rates. 2020PhRvL.125c2701F recently measured the ratio of to capture using superconducting tunnel junctions. The recommendation of both SF I and SF II was

### v2

@@ VI.3 Experimental progress on
Since 2011RvMP...83..195A, five new experiments have been reported: Four from the Madrid and ATOMKI groups by the activation technique, at relatively high center of mass energies [2012PhRvC..86c2801C, 2013NuPhA.908....1B, 2019PhRvC..99e5804S, 2023PhRvC.108b5802T], and one from the Notre Dame group using the prompt- method [2013PhRvC..87f5804K].
@@
Following the approach adopted in 2011RvMP...83..195A, we again only use the recoil and activation data for the fits below. The -ray angular distribution is not known experimentally [2007PhRvC..76e5801B], and the resultant uncertainty increases the common-mode error for the prompt- studies somewhat. (An experiment to address previously-raised concerns about angular distribution [2007PhRvC..76e5801B] has recently concluded at Felsenkeller Dresden. The data suggest a higher than expected anisotropy but are so far only available in the form of a PhD thesis [Turkat23-PhD]. If confirmed, they may lead to corrections of a few percent for some of the in-beam experiments where only one angle was instrumented.) In addition, all of the “modern” works reporting prompt- data except for Notre Dame [2013PhRvC..87f5804K] also include data obtained with other methods, leading to partial correlations between data sets that would complicate fitting.
@@
Following this restriction, our data selection proceeds as follows. First, the four data sets previously used in 2011RvMP...83..195A are carried over here: Weizmann [2004PhRvL..93z2503S], LUNA [2006PhRvL..97l2502B, 2007PhRvC..75f5803C, 2007PhRvC..75c5805G] (only the activation data), Seattle [2007PhRvC..76e5801B] (only the activation data), and ERNA [2009PhRvL.102w2502D].
@@
Two of the archival data sets excluded by the 1998 cutoff date in 2011RvMP...83..195A merit some further discussion here: The data of 1982PhRvL..48.1664O, 1984NuPhA.419..115O consist of two points measured by activation in a 3He gas cell. However, their uncertainties are not separated into statistical and systematic components. Including them would have required an uncertain guess on how to divide the error bars, and the fit result would barely change, so we left them out. The work by 1983PhRvC..27...11R consists of one data point that was obtained by averaging two separate activation measurements with a 3He and a 4He gas cell, respectively. We do not use this result because we lack details to verify the background subtraction, an issue that was raised during our meeting by one of its authors. We also choose not to use the activation study by 1983ZPhyA.310...91V. In that experiment there were thick entrance foils and the 4He beam was completely stopped inside a high-pressure 3He gas cell. This gave an integrated measurement over a wide energy range, so that the analysis depends strongly on the assumed shape of the curve, and the result was reported only as an extrapolated .
@@
We now consider the new data since 2011RvMP...83..195A. An activation experiment in Madrid [2012PhRvC..86c2801C] used a 3He beam incident on a 4He gas cell, reporting three data points. Another activation experiment was reported by the ATOMKI group, using a 4He beam on 3He gas cells [2013NuPhA.908....1B]. Two higher-energy campaigns at ATOMKI were again performed using the activation method. These latter data are at = 2.5-4.4 MeV [2019PhRvC..99e5804S] and = 4.3-8.3 MeV [2023PhRvC.108b5802T], respectively, above the energy range suitable for halo EFT, and therefore not included.
@@
A detailed study by the Notre Dame group reported 17 data points [2013PhRvC..87f5804K], using the primary -ray from ground state capture and the secondary -ray from the deexcitation of the 429 keV first excited state of 7Be. No activation data are reported in that work, and the -ray detector was placed at just one angle, 90∘. Since only prompt- data are reported, we did not include this data set. In order to test the effects of this decision, we repeated some of our fits (see below, Section VI.4) with the modern prompt- experiments included: the 17 points from 2013PhRvC..87f5804K, 3 prompt- points from LUNA [2007PhRvC..75f5803C] and 8 from Seattle [2007PhRvC..76e5801B]. The extrapolated S-factor changed by less than 1%, well within the error bars for the recommended .
@@
Finally, an indirect experiment using the 6Li(3He,d)7Be reaction and the ANC technique has recently been reported [2020PhLB..80735606K]. It is left out of our fits due to the additional normalization and theory uncertainties involved in determining an ANC from a transfer experiment, which are larger than the errors in the S-factor measurements used for the present fits. All the data used in the present fits are summarized in Table 5.
@@
As in 2011RvMP...83..195A, we model data uncertainties as consisting of a component that is independent for each point and a common-mode component that applies to all data from a given experiment as a multiplicative factor. This separation is well-documented for all of the modern data and is also shown in Table 5. Except at the lowest energies, the common-mode error typically dominates.
@@ VI.4 Data fitting
We carried out both frequentist fits that minimize and Bayesian fits based on the posterior probability distribution of a likelihood function computed from . In the Bayesian fits parameters, extrapolated , and their errors were determined by Monte Carlo sampling of the parameter space. The relatively large number of model parameters and the significant fitting degeneracies between some of them make Bayesian analysis a natural choice for finding best fit and confidence intervals for the multi-parameter models applied here. Two of the groups ran into serious difficulty with the frequentist fits due to parameter degeneracies and shallow local minima; during this study they were unable to produce frequentist fits in which they had confidence.
@@
One set of fits for the 2018EPJA...54...89H halo EFT was produced by minimization, proceeding in two steps: first LO parameters were fitted to keV data, and then the results were taken as initial values in the search to minimize over all parameters in the NLO theory for the full set of data. These fits gave keV b with the scattering constraint and keV b without (uncertainties being propagated in linear approximation using covariances and partial derivatives). Formally the EFT truncation error from stopping at NLO corresponds to an additional theoretical error that can shift S-factors by 10%. However, any fitted curve is constrained by low-energy data, which in some sense become effectively renormalization conditions of the field theory; the error on extrapolated should probably be smaller than 10% by an amount that is hard to estimate.
@@ VI.5 Recommended value
Because the fit to the NLO EFT of 2018EPJA...54...89H gives the central result among those attempted here, we adopt it for our recommended . For ease of adoption, it is noted that this recommended curve can be empirically parameterized using the same shape as in 2011RvMP...83..195A, by the numerical equation (following a customary form in the past literature [1987ApJ...319..531K, 2011RvMP...83..195A])
@@
Propagation of uncertainties from measurement to extrapolated is relatively straightforward and unambiguous (at least in the Bayesian fitting). However, the uncertainty due to theoretical approximation is more complicated, and error estimation on extrapolated quantities in EFTs remains an open area of research; estimates from the groups working on are discussed above. We estimate the overall theoretical uncertainty on our recommended S-factor to be the measured by the 4% dispersion among fitted curves; this is comparable to the more formal estimates.
Our recommendations for the zero energy astrophysical S-factor and the low-energy slope are then
@@
The first error bar for is given by the fit uncertainty, conservatively using the largest value among the various fits, and it encodes the experimental uncertainty. The second error bar reflects the 4% dispersion among theoretical formalisms. The two errors can be added in quadrature for applications. Our recommendation is essentially the same as in 2011RvMP...83..195A, but derived from completely different theoretical approaches that are much more directly connected to the data and much less dependent on model assumptions. It is also based on data over a wider energy range. The dispersion of and among models at the solar Gamow peak is illustrated in Table 6, along with some indication of how fitted scattering lengths and squared ANCs vary among frameworks. (The latter are highly degenerate with short-range parameters in the fitting; in the 2018EPJA...54...89H EFT they are sensitive to small variations in the -wave effective ranges that are fitted instead of ANCs.)
@@ VI.6 Recommendations for future work on
Our third recommendation is a study of the astrophysical S-factor over a wide energy range, 50-1500 keV and extending to even lower energies when technically feasible, with small point-to-point errors. Even with a large common-mode error, this would significantly constrain theoretical curve shapes near threshold, and providing an independent check on the influence of scattering data on .
@@
The hep reaction ()
The reaction , also known as the hep reaction, is one of the possible processes taking place after the , and it produces the most energetic neutrinos, with an endpoint energy of 18.8 MeV. It is however the least probable, with its S-factor, and consequently its rate, even beyond the reach of current experiments. This is due to the fact that the hep reaction is induced by the weak interaction and further suppressed by Coulomb barrier and by the fact that the leading one-body axial (Gamow-Teller) current operator cannot connect the main -state components of the and initial- and final-state wave functions. As a consequence, the one-body axial current non-zero contribution is due to the small components of the and wave functions. Furthermore, other contributions, such as weak magnetism and other one-body corrections to the vector current, and two-body vector and axial currents, and the -waves in the initial state, normally suppressed at solar temperatures, are significantly enhanced. This is further complicated by cancellations between the one-body and two-body axial current contributions. As a consequence, similarly to the pp fusion, the hep reaction cross section is too small to be measured in laboratory and only theoretical predictions are available.
The most recent studies of this reaction, already evaluated in 2011RvMP...83..195A, are those of 2000PhRvC..63a5801M and 2003PhRvC..67e5206P. The first one was performed within a phenomenological approach, similar to the calculation of 1998PhRvC..58.1263S for the pp fusion. In this particular case, the and initial and final nuclear wave functions were obtained with the correlated Hyperspherical Harmonics (CHH) variational method [1995FBS....18...25V, 1998PhRvL..81.1580V], using the Argonne (AV18) two-nucleon potential [1995PhRvC..51...38W] augmented by the Urbana IX (UIX) three-nucleon interaction [1995PhRvL..74.4396P]. In this way, the binding energies of 3He and 4He, and the singlet and triplet scattering lengths were in good agreement with experiment. The weak vector and axial transition operators were obtained within a phenomonological approach. In particular, the largest two-body contribution to the axial current arising from the excitation of intermediate -isobar degrees of freedom was included using the transition-correlation operator scheme of 1991PhRvC..44..619C, 1992PhRvC..45.2628S and fixing the nucleon-to- axial coupling constant to reproduce as in the pp case [1998PhRvC..58.1263S].
@@
Following the results of 2000PhRvC..63a5801M and 2003PhRvC..67e5206P, 2011RvMP...83..195A recommended for the zero-energy S-factor, , the value keV b. Three recommendations were also made: (i) to perform new studies with a broad spectrum of Hamiltonian models, in order to properly access the theoretical uncertainty; (ii) to study other weak reactions for which experiments can test the predictions made by employing the same theoretical ingredients as those used for the hep reaction calculation; (iii) to further understand the relation between the hep reaction and the so-called hen process (). Of these recommended works, only the second one has been performed so far, applying the theoretical framework used for the hep study to the muon capture reactions on light nuclei [2011PhRvC..83a4002M].
Given the lack of further studies on the hep reaction after 2011RvMP...83..195A, we decide to maintain
as recommended value. Notice that we could have increased by %, in analogy with , due to the updated values of input parameters (mostly ). However, we have decided not to apply such an increase, because it does not derive from a direct calculation, and, most of all, because such an increase lies well within the quoted % uncertainty. However, given the advances in ab initio studies of the last decade and in the development of the EFT framework, we update the recommendations of 2011RvMP...83..195A as follows: (i) the hep process should be revisited within a fully consistent EFT approach, similarly to what has been done in these years for the pp fusion and other processes, as, for instance, muon captures on light nuclei [2012PhRvL.108e2502M, 2018PhRvL.121d9901M, 2018PhRvC..98f5506A, 2023FrP....1049919C, 2023PhRvC.107f5502B, PhysRevC.109.035502]; (ii) the initial and final wave functions should be calculated with the most accurate ab initio methods, as the (uncorrelated) HH method [2008JPhG...35f3101K, 2020FrP.....8...69M], which has been proven to provide more accurate bound and scattering wave functions than the CHH method, using both local and non-local interactions; (iii) the relation between the hep and the hen process should be understood. The hen process has been studied within EFT and the HH method by 2022PhRvC.105a4001V, but a consistent parallel study of both reactions is still missing.
@@
Electron capture by p+p and
Electron capture reactions are the sources of lines found in the solar neutrino spectrum. In particular, electron capture on , i.e., the process d, also known as the pep reaction, competes with fusion and depends on the same nuclear matrix element. Therefore, the ratio between the pp and pep rates is independent of nuclear physics. Based on this consideration, in 2011RvMP...83..195A the result of 1998RvMP...70.1265A, which is based on the work of 1969ApJ...155..501B, was multiplied by the radiative corrections calculated by 2003PhRvC..67c5502K, leading to the final result given by
where is the density in units of g/cm3, is the mean molecular weight per free electron, and is the temperature in units of MK. The range of validity is . Given that no new evaluation of the pep rate has been performed since 2011RvMP...83..195A, the result of Eq. (20) represents also the present recommended value.
Competition between electron and proton capture on 7Be fixes the branching ratio of the ppII and ppIII chains, and thereby the 7Be and 8B neutrino fluxes. Superallowed electron capture on the ground state of 7Be leads to the ground state or the , 478 keV first excited state of 7Li with a measured (terrestrial) branching ratio of 10.44(4)% [2002NuPhA.708....3T]. It is customary to calculate the solar decay rate in terms of the measured terrestrial decay rate, which has the advantage of removing the nuclear physics dependence, but it requires a proper calculation of the densities of electrons at the nucleus. This has usually been done by considering continuum and bound electrons separately, using the Debye-Hückel (DH) approximation pioneered by 1954AuJPh...7..373S for the screening of bound electrons in the solar plasma; the capture rate of continuum electrons is not significantly influenced by electron screening. Furthermore, it is necessary to calculate the atomic probability densities governing the and terrestrial electron capture rates. 2020PhRvL.125c2701F recently measured the ratio of to capture using superconducting tunnel junctions. The recommendation of both 1998RvMP...70.1265A and 2011RvMP...83..195A was